# Data-Efficient Sleep Staging with Synthetic Time Series Pretraining

### v1

@@ Introduction
Efforts to mitigate the scarcity of large datasets have primarily followed two paths: (1) the development of network architectures that incorporate constraints mirroring the data’s intrinsic characteristics, such as symmetries [5], and (2) enhancing model performance through the use of additional or cross-domain data to learn effective priors. Pertaining to the first path, a common feature in time series processing networks is the use of convolutional layers. These layers are designed to be translation-equivariant [6], which ensures that a temporal shift in the input only affects the output by the same shift. This characteristic enables consistent network responses to temporal patterns, regardless of their temporal location, while reducing the number of model parameters compared to architectures lacking such constraints. For the second path, a variety of strategies have been proposed to learn useful priors from data. One approach is data augmentation, in which time series are transformed while preserving their annotations (labels) to artificially expand the dataset [7, 8]. Deep neural networks trained on such augmented datasets implicitly learn to become invariant under these transformations, which can lead to better out-of-sample prediction performance. Another strategy is transfer learning [9], a two-step process in which neural networks are trained on one task using a large dataset (pretraining step [10]) and then adapted to learn the actual task of interest using another (usually much smaller) dataset (fine-tuning step). A variant of this idea is self-supervised learning [11, 12], which allows neural networks to be pretrained on large and heterogeneous datasets without explicitly labeled examples. Finally, generative models such as GANs can be used to sample new time series to extend existing datasets [13, 14]. Such generative models approximate a data distribution and require large heterogeneous datasets for training. While all of these approaches have been demonstrated to be able to improve the performance of neural networks, they still rely on large empirical datasets for training.
@@
Inspired by these advances, we hypothesize that pretraining exclusively on synthetic time series data generated from simple random processes can also yield effective priors for sleep staging. Given the importance of frequencies for sleep stage scoring and other EEG-based applications [17, 18], we introduce a pretraining method that centers on generating synthetic time series data with specific frequency content (see Fig. 1). During pretraining, deep neural networks learn to accurately predict the frequencies present in these synthetic time series. We observe that this conceptually simple pretraining task, which we call “frequency pretraining” (FPT), allows a deep neural network to detect sleep stages with better accuracy compared to fully supervised training when few samples (few-samples regime) or data from few subjects (few-subject regime) are available for fine-tuning (see Fig. 2). The success of our method underscores the essential role of frequency content in enabling neural networks to accurately and reliably discern sleep stages. We consider pretraining techniques leveraging synthetic data, like the one we propose, as a promising area of research, offering the potential to develop models in sleep medicine and neuroscience that are particularly suited for scenarios involving small datasets. To facilitate testing and further advancements, we make the source code of our method publicly available [19].

### v2

@@ 1 Introduction
Efforts to mitigate the scarcity of large datasets have primarily followed two paths: (1) the development of network architectures that incorporate constraints mirroring the data’s intrinsic characteristics, such as symmetries [7], and (2) enhancing model performance with additional or cross-domain data to learn effective priors. Pertaining to the first path, a common feature in time series processing networks is the use of convolutional layers. These layers are designed to be translation-equivariant [18], which ensures that a temporal shift in the input only affects the output by the same shift. This characteristic enables consistent network responses to temporal patterns, regardless of their temporal location, while reducing the number of model parameters compared to architectures lacking such constraints. For the second path, a variety of strategies have been proposed to learn useful priors from data. One approach is data augmentation, in which time series are transformed while preserving their annotations (labels) to artificially expand the dataset [31, 23]. Deep neural networks trained on such augmented datasets implicitly learn to become invariant under these transformations, which can lead to better out-of-sample prediction performance. Another strategy is transfer learning [12], a two-step process in which neural networks are trained on one task using a large dataset (pretraining step [37]) and then adapted to learn the actual task of interest using another (usually much smaller) dataset (fine-tuning step). A variant of this idea is self-supervised learning [32, 2], which allows neural networks to be pretrained on large and heterogeneous datasets without explicitly labeled examples. Finally, generative models such as VAEs, GANs, and diffusion models can be used to sample new time series to extend existing datasets [22, 8, 46]. Such generative models approximate a data distribution and require large heterogeneous datasets for training. While all of these approaches have been demonstrated to be able to improve the performance of neural networks, they still rely on large empirical datasets for training.
@@
Inspired by these advances, we hypothesize that pretraining exclusively on synthetic time series data generated from simple random processes can also yield effective priors for sleep staging. Given the importance of frequencies for sleep stage scoring and other EEG-based applications [5, 33], we introduce a pretraining method called “frequency pretraining” (FPT) that centers on generating synthetic time series data with specific frequency content. During pretraining, deep neural networks learn to accurately predict the frequencies present in these synthetic time series. Despite the deliberate simplicity of our synthetic data generation process and the inherent domain shift between synthetic and EEG data, we observe that FPT allows a deep neural network to detect sleep stages more accurately than fully supervised training when few samples (few-sample regime) or data from few subjects (few-subject regime) are available for fine-tuning. The success of our method underscores the essential role of frequency content in enabling neural networks to accurately and reliably discern sleep stages. We consider pretraining techniques leveraging synthetic data, like the one we propose, as a promising area of research, offering the potential to develop models in sleep medicine and neuroscience that are particularly suited for scenarios involving small datasets. To facilitate testing and further advancements, we make the source code of our method publicly available [19].
Contributions
  • Novel synthetic pretraining approach: We introduce “frequency pretraining” (FPT), a pretraining approach using synthetic time series with random frequency content, eliminating the need for empirical EEG data during pretraining.
  • Demonstrated data efficiency: We demonstrate superior sleep staging performance of FPT in few-sample and few-subject regimes across three datasets, with comparable results to fully supervised methods when data is abundant.
  • Analysis of frequency-based priors: We evaluate the role of frequency information in our pretraining task and how synthetic sample diversity affects fine-tuning.
  • Comparison with self-supervised methods: We benchmark our approach against established self-supervised learning methods, demonstrating that synthetic data pretraining can achieve comparable results without requiring EEG data for pretraining.
2 Results
Our approach, illustrated in Figure 1 and detailed in Section 4, is based on a two-phase training process that combines pretraining on synthetic time series with fine-tuning on clinical sleep data. During the pretraining phase, we generate synthetic time series composed of sine waves with random frequencies drawn from predefined frequency ranges (frequency bins). These synthetic signals are then used to train a deep neural network whose feature extractor, $f$, learns to extract useful features, while the classifier, $c_{p}$, learns to predict the frequency bins from which the frequencies were drawn to generate the synthetic time series. After pretraining, the feature extractor is transferred to the fine-tuning phase, where it is applied to real EEG and electrooculography (EOG) data to classify sleep stages. In this phase, the model processes sequences of eleven consecutive sleep epochs, indexed i5i-5 to $i+5$, with the feature extractor producing features $h_{i}$ from each epoch. Another classifier $c_{f}$ then aggregates these features to predict the sleep stage (Wake, N1, N2, N3, REM) for the central sleep epoch $i$.
We evaluated our approach on three publicly available datasets, DODO/H, Sleep-EDFx, and ISRUC, which provided data from 276 subjects, including both healthy individuals and those with various medical conditions (see Section 4.2.1). For sleep staging performance, we tracked the Macro-F1 score, with higher values indicating better classification accuracy across sleep stages. During the pretraining phase, we assessed the model’s ability to predict frequency bins using the Hamming metric and the accuracy (see Section 4).
2.1 Training Configurations
We compared the performance of pretrained models against the performance of non-pretrained models in scenarios with varying amounts of training data. In particular, we studied the performance of our approach in few-sample and few-subject regimes, where the greatest benefit was expected. Furthermore, we analyzed the priors that the model learned during pretraining and the role of frequency information in the learned features. Finally, we investigated whether these features could be further improved by fine-tuning the feature extractor. To enable these investigations, we created four training configurations.